OpenStudy (anonymous):

To measure the speed of a current, scientists place a paddle wheel in the stream and observe the rate at which it rotates. If the paddle wheel has radius 0.20m and rotates at 100rpm, find the speed of the current in m/s.

OpenStudy (anonymous):

1 rpm = 2pi rad/min = 2pi rad / 60s = 1 * (pi/30) rad/s 100 rpm = 100 * (pi/30) rad/s = w R = 0.20m V = w R V = 100 * (pi/30) * 0.20

OpenStudy (anonymous):

100 times circumference per minute.....

OpenStudy (anonymous):

V = 100 * (pi/30) * 0.20 V = 2.0944 m/s

OpenStudy (anonymous):

okay thanks